OpenStudy (anonymous):

Use the substitution method to solve the system of equations. Choose the correct ordered pair. 2x + 3y = 29 x = 4 A. (4, 7) B. (4, 11) C. (7, 4) D. (3, 2) could you maybe walk me through this just to make sure i got it @Zale101

OpenStudy (zale101):

you already have x=4, a variable is already given, s0 plug x as 4 to the equation to get y

OpenStudy (zale101):

plug it to the first equation, i think you can handle this xD Like you did in the first problem because this one is simply easier

OpenStudy (anonymous):

A? i got 7

OpenStudy (zale101):

how did you get 7?

OpenStudy (anonymous):

2x4+3y=29

OpenStudy (zale101):

2(4)+3y=29 8+3y=29 3y=21 y=7

OpenStudy (anonymous):

Ok! So then it would be (4,7)

OpenStudy (anonymous):

Correct?